the larger of two numbers is less than six times the smaller. when the smaller number is subtracted from the larger, the difference is 13. find the two numbers
Answer by ankor@dixie-net.com(22740)

the larger of two numbers is less than six times the smaller.
a < 6b
when the smaller number is subtracted from the larger, the difference is 13.
a - b = 13
a = (b+13)
:
find the two numbers
In the first equation, replace a with (b+13)
b + 13 < 6b
13 < 6b - b
13 < 5b
divide both sides by 5
2.6 < b
Lets say b = 3
then
a = 3 + 13
a = 16
:
a = 16, b = 3
:
Check in the first equation
16 < 6(3)